please answer both
3. Differentiate the following using the Difference Quotient. a. f(x) 3x b. g(x) 5 cos(x)
Answer #1

a)

we have

f(x)=3\sqrt{x}

so,

f(x+h)=3\sqrt{x+h}

now,

(x)lim+h)-

f'(x)=\lim_{h \to 0}\frac{3\sqrt{x+h}-3\sqrt{x}}{h}

r(z) = 3 linn

3 lim

f(r) 3 lim

f(r) 3 lim

f(x) =-

f ()3 2

b)

we have

g(x)=5cos(x)

so,

g h)5cos( h) 5(cos()cos(h) - sin(x)sin(h))

now,

g'(x)=\lim_{h \to 0}\frac{g(x+h)-g(x)}{h}

g'(x)=\lim_{h \to 0}\frac{5(cos(x)cos(h)-sin(x)sin(h))-5cos(x)}{h}

g'(x)=5\lim_{h \to 0}\frac{cos(x)(cos(h)-1)-sin(x)sin(h)}{h}

g'(x)=5\left [cos(x)\lim_{h \to 0}\frac{cos(h)-1}{h}-sin(x)\lim_{h \to 0}\frac{sin(h)}{h} \right ]

we know that,

\lim_{x \to 0}\frac{cos(x)-1}{x}=0 \ and \ \lim_{x \to 0}\frac{sin(x)}{x}=1

we can say that

5 cos(r)(0)- sin x)

g'(x)=-5sin(x)